OpenStudy (anonymous):

A bag of ten tennis balls that are three different colors. Chaz picks a ball from the bag 40 times without looking and returns it to the bag each time. The results of his experiment are 4 pink, 21 yellow, and 15 purple. How many of each color ball do you think are in the bag? Explain your answers.

OpenStudy (whpalmer4):

We had 40 trials. Number of each color that came out / 40 trials gives us an approximation of the fraction of the whole that color represents. Then multiply that fraction by the number of balls in the bag to get the number of balls of each color. For example, 4 pink balls were drawn, so 4/40 = 1/10 is the fraction that is pink. 1/10 * 10 = 1 pink ball. When you are done, the sum of the numbers of balls of each color should add up to the number of balls in the bag.